OpenStudy (bloomlocke367):

Write an equation of a circle with the given center and radius. center (3, 4) and radius 6.

OpenStudy (bloomlocke367):

@campbell_st

OpenStudy (bloomlocke367):

it's been awhile since I've done Conics.... sorry cx

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2 = r\] where the center is (h,k) and radius = r.

OpenStudy (bloomlocke367):

ohhh okay

OpenStudy (bloomlocke367):

thank you

OpenStudy (anonymous):

Np

OpenStudy (bloomlocke367):

so \((x-3)^2+(y-4)^2=6?\)

OpenStudy (anonymous):

That should be r^2, so it's 6^2

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2 = r^2\]

OpenStudy (bloomlocke367):

oooo so \((x-3)^2+(y-4)^2=36\)

OpenStudy (anonymous):

Yeah :)

OpenStudy (bloomlocke367):

thank you c: can you help with more, maybe?